OpenStudy (anonymous):

square root of 5 x 4= 4 x _. illustrate the commutative property.

OpenStudy (angela210793):

\[\sqrt{5x*4} ? or \sqrt{5*4}?\]

OpenStudy (angela210793):

if it is the 1st:= 2sqr(5x) id it is the 2nd:=2sqr5... hope ur looking 4 this :)

OpenStudy (toxicsugar22):

hi angela can u help me

OpenStudy (angela210793):

Idk maybe...

OpenStudy (toxicsugar22):

4x^2-6y^2+32x-24y+16=0 give the exact coordinated of an four points

OpenStudy (toxicsugar22):

or anyone help me

OpenStudy (toxicsugar22):

4x^2-6y^2+32x-24y+16=0 give the exact coordinates of any four points on the hyperbole

OpenStudy (toxicsugar22):

ok then nobody will help me

OpenStudy (angela210793):

let me see some formulas 1st pls...i don't remember them well

OpenStudy (toxicsugar22):

ok well then i am going to go

OpenStudy (toxicsugar22):

so i will see u later

OpenStudy (angela210793):

(4X^2 +32X +64)-64=(X+4)^2 - 64....do the same for y....and u'll get the right eq. of hiperbole.... c^2-a^2=b^2,hiperbola intercepts (Ox) at (a;0) and (-a;0) I'm really late 4 school and i haevn't had breakfast yet...I'll see it later :) I'm sorry